OpenStudy (anonymous):

Casey has a small business making dessert baskets. She estimates that her fixed weekly costs for rent, electricity, and salaries are $200. The ingredients for one dessert basket cost $2.50. Casey’s total costs for last week were $500. How many dessert baskets did she make? a. 300 b. 12 c. 120 d. 280

OpenStudy (anonymous):

now fromtotal costs we take out the weekly costs 500-200=300 now from 300 she can make 120 dessert baskets 300/2.50 =120

OpenStudy (texaschic101):

200 + 2.50x = 500 now solve for x